OpenStudy (anonymous):

graph the function f(x)=x^2+4x-12 and determine the number of zeros it has

OpenStudy (anonymous):

make f(x)=0 and you'll find the zero(s)

OpenStudy (anonymous):

can u show me how i dnt get it

OpenStudy (anonymous):

Factor x^2+4x-12

OpenStudy (anonymous):

how

OpenStudy (anonymous):

(x-2)(x+6)=0 so x-2=0 and x+6=0 x=2 and x=-6

OpenStudy (anonymous):

okay but how do i find the zeros

OpenStudy (anonymous):

Those are the zeros.

OpenStudy (anonymous):

oh thANX